OpenStudy (anonymous):

Graph the two equations to solve the system. What is the solution? y = 2x - 4 and 2x + 3y = -12 A. (0, -4) B. (4, 4) C. no solution D. infinitely many solutions

hero (hero):

You can graph it, or you can just solve it algebraically: y = 2x - 4 y = (-12 - 2x)/3 y = y 3(2x-4) = -12 - 2x 6x - 12 = -12 - 2x 8x = 0 x = 0/8 x = 0 y = -4

OpenStudy (anonymous):

(0,-4) is d solution

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

Plot attached.